OpenStudy (anonymous):

GIVING medals!!!!!! Use the rules for integer addition to simplify the expression. Positive twelve plus negative eleven equals what value? negative twenty-three negative one positive one positive twenty-three --------------------------------------------------------------------------------

sammixboo (sammixboo):

When you are adding integers with a positive you always switch the addition sign with a subtraction sign and then subract the numbers, and then take the number with the greater absolute value and take the integer sign from that and plop it onto the answer. 12 + (-11) = 12 - (-11) = Since 12 has the greater absolute value, the answer will be positive.. SO your answer is +1